OpenStudy (anonymous):

simplify (49a^2-36b^2)/(49a-42b)

jimthompson5910 (jim_thompson5910):

Can you factor 49a^2-36b^2 ?

OpenStudy (anonymous):

yeah (7a-6b)(7a-6b)

jimthompson5910 (jim_thompson5910):

good

jimthompson5910 (jim_thompson5910):

what about 49a-42b ?

jimthompson5910 (jim_thompson5910):

oh wait, I misread

jimthompson5910 (jim_thompson5910):

it should be (7a-6b)(7a+6b)

jimthompson5910 (jim_thompson5910):

since it's a difference of squares

OpenStudy (anonymous):

so (7a-6b)(7a+6b)/(7a-6b)

jimthompson5910 (jim_thompson5910):

close

jimthompson5910 (jim_thompson5910):

the denominator should be 7(7a-6b)

jimthompson5910 (jim_thompson5910):

because you factor out 7

jimthompson5910 (jim_thompson5910):

So.... \[\Large \frac{49a^2-36b^2}{49a-42b}\] \[\Large \frac{(7a-6b)(7a+6b)}{49a-42b}\] \[\Large \frac{(7a-6b)(7a+6b)}{7(7a-6b)}\] \[\Large \frac{7a+6b}{7}\]

OpenStudy (anonymous):

yeah i see it now

OpenStudy (anonymous):

thank you!

jimthompson5910 (jim_thompson5910):

you're welcome